OpenStudy (anonymous):

Rational expressions: (1)/(r²-1) = (2)/(r²+r-2) These are fractions !

OpenStudy (saifoo.khan):

r^2 - r

OpenStudy (anonymous):

r equals 0

OpenStudy (anonymous):

\[1/(r-1)(r+1)=2/(r+2)(r-1)\] \[1/(r+1)=2/(r+2)\]

OpenStudy (anonymous):

2(r+1)=r+2 2r+2=r+2 r=0

OpenStudy (anonymous):

you have to add that: r is NOT = -2 r is not = -1 r is not = 1 (can't divide by 0)

OpenStudy (saifoo.khan):

1 0r 0

OpenStudy (anonymous):

Can't be 1, zero divisor

OpenStudy (saifoo.khan):

it a quardatic eq.

OpenStudy (anonymous):

As presented, cannot be 1.